# Two-color partitions and overpartitions: a combinatorial proof

### v1

@@ Abstract.
George Andrews and Mohamed El Bachraoui recently explored identities for two-color partitions [2]. In particular, they studied the connection between two-colored partitions and overpartitions. Their proofs were analytical, but they conjectured combinatorial proofs of their results. In this paper we use two-modular diagrams to give a combinatorial proof of their main result.
@@ 1. Introduction
Throughout the paper we let $n$ denote a nonnegative integer. Recall that a partition of $n$ is a tuple $(\lambda_{1},\dots,\lambda_{k})$ such that $\lambda_{1}+\dots+\lambda_{k}=n$ and $\lambda_{1}\geq\lambda_{2}\geq\dots\geq\lambda_{k}\geq 1$. Let us call a partition two-colored if each summand is colored in either green or blue. For example, the tuple $(9_{g},5_{b},4_{g},2_{g})$ is a partition of 20.
@@
Recently George E. Andrews and Mohamed El Bachraoui [2] established the following connection between two-colored partitions and overpartition.
@@
Their proof is analytical; however, they conjectured combinatorial proof of Theorem 1.2. In particular, they conjectured
@@
(Section 9, [2]). It would be interesting to give combinatorial proofs for the identities in Theorem 1.2.
@@ 2. Combinatorial proof of item a)
Recall that Euler’s theorem [1] states that the number of partitions of a nonnegative integer $n$ into odd parts coincides with the number of partitions of $n$ into distinct parts. This theorem has a well-known combinatorial proof.
@@
We point out that D. Chen and Z. Zou [3] also gave a combinatorial proof of item a).
@@ 3. Two-modular diagrams
Let us introduce the two-modular diagram of $\lambda$. In these diagrams $\lambda_{g}$ parts are represented in rows above the main diagonal and $\lambda_{b}$ parts are represented in columns below the main diagonal ($\lambda_{e}$ parts are not represented in the diagram). Each odd part is represented in the diagram the following way: if a part equals $2k+1$, then its representation would have $k$ squares and 1 triangle. In other words, a square in the diagram represents number “two” and a triangle represents number “one”.
@@ 4. Combinatorial proof of items b) to e)
Thus if we ignore two aforementioned partitions, we establish a one-to-one correspondence between $E_{0}(n)\leftrightarrow E_{1}(n)$ and $E_{2}(n)\leftrightarrow E_{3}(n)$.
@@
It deserves to be mentioned, that the same idea of an almost-involution happens to be in the proof of Euler’s pentagonal theorem [1]. Moreover, the generalization of two-modular diagram allows to prove combinatorically the Jacobi triple product identity [4].

### v2

@@ Abstract.
George Andrews and Mohamed El Bachraoui recently explored identities for two-color partitions [andrews2024twocolorpartitionsoddsmallest]. In particular, they studied the connection between two-colored partitions and overpartitions. Their proofs were analytical, but they also conjectured combinatorial proofs of their results. In this paper we use two-modular diagrams to give a combinatorial proof of their main result.
@@ 1. Introduction
Throughout the paper we let $n$ denote a nonnegative integer. Recall that a partition of $n$ is a tuple $(\lambda_{1},\dots,\lambda_{k})$ such that $\lambda_{1}+\dots+\lambda_{k}=n$ and $\lambda_{1}\geq\lambda_{2}\geq\dots\geq\lambda_{k}\geq 1$. Let us call a partition two-colored if each summand is colored in either green or blue. For example, the tuple $(9_{g},5_{b},4_{g},2_{g},2_{b},2_{b})$ is a partition of 24.
@@
Recently George E. Andrews and Mohamed El Bachraoui [andrews2024twocolorpartitionsoddsmallest] established the following connection between two-colored partitions and overpartitions.
@@
Their proof is analytical; however, they conjectured a combinatorial proof of Theorem 1.2. In particular, they conjectured
@@
(Section 9, [andrews2024twocolorpartitionsoddsmallest]). It would be interesting to give combinatorial proofs for the identities in Theorem 1.2.
@@ 2. Combinatorial proof of item a)
Recall that Euler’s theorem [andrews] states that the number of partitions of a nonnegative integer $n$ into odd parts coincides with the number of partitions of $n$ into distinct parts. This theorem has a well-known combinatorial proof.
@@
We point out that D. Chen and Z. Zou [articlecombinatoricalproofs] also gave a combinatorial proof of item a).
@@ 3. Two-modular diagrams
Let us introduce the two-modular diagram of $\lambda$. In these diagrams $\lambda_{g}$ parts are represented in rows above the main diagonal and $\lambda_{b}$ parts are represented in columns below the main diagonal ($\lambda_{e}$ parts are not represented in the diagram). Each odd part is represented in the diagram the following way: if a part equals $2k+1$, then its representation would have $k$ squares and 1 triangle. In other words, a square in the diagram represents the number “two” and a triangle represents the number “one”.
@@ 4. Combinatorial proof of items b) to e)
Thus if we ignore the two aforementioned partitions, we establish a one-to-one correspondence between $E_{0}(n)\leftrightarrow E_{1}(n)$ and $E_{2}(n)\leftrightarrow E_{3}(n)$.
@@
It deserves to be mentioned, that the same idea of an almost-involution happens to be in the proof of Euler’s pentagonal theorem [andrews]. Moreover, the generalization of two-modular diagram allows one to prove combinatorically the Jacobi triple product identity [jacobi].
@@ 5. Acknowledgments
This work was performed at the Saint Petersburg Leonhard Euler International Mathematical Institute and supported by the Ministry of Science and Higher Education of the Russian Federation (agreement no. 075-15-2025-343).
6. Data availability statement
The authors declare that the data supporting the findings of this study are available within the paper.